OpenStudy (anonymous):

The sum of two consecutive positive odd integers is between 7 and 17. Which pair of integers does not solve this problem? A. 3 and 5 B. 5 and 7 C. 1 and 5

Parth (parthkohli):

3 + 5 = 8 \(\checkmark\) 5 + 7 = 12 \(\checkmark\) 1 + 5 = 6 \(\text{Nope}\)